# On the Expressive Power of a Variant of the Looped Transformer

### v1

@@ 5.2 Impact of Hyperparameters
In this subsection, we conduct the empirical analysis of the impact of the hyperparameters on the AlgoFormer. The visualized results are provided in Appendix B.2. Notably, we observe that when training on higher values of $T$ and $\Delta T$, the AlgoFormer achieves stability with prolonged inference iterations. This observation suggests that the designed transformer exactly performs specific algorithms during its operation. However, it is important to note that larger values for $T$ and $\Delta T$ correspond to deeper transformers, leading to increased computational costs during both training and inference. On the contrary, smaller values for $T$ and $\Delta T$ may result in suboptimal performance. Striking a balance between the desired expressiveness and computational efficiency becomes imperative, highlighting the existence of a trade-off in these hyperparameters. Additionally, we investigate the influence of the number of heads and layers in the looped transformer $\text{TF}_{\text{loop}}$. We observe that a greater number of heads and layers contribute to the improved expressiveness, as shown in Figure 5. However, the overly many layers and heads may diminish the training and generalization.
5.3 Transformers May be Smarter than Human-Designed Algorithms
In this subsection, we leverage the AlgoFormer to address linear regression problems, presenting it as a pedagogical example for comparison with traditional methods such as gradient descent and Newton’s method. The hyperparameters in these experiments are detailed in Appendix B.3, with their selection grounded in a comprehensive grid search. According to the visualized results shown in Figure 6, transformers with fewer iterations outperform gradient descent and perform comparably and even better than Newton’s method. This observation suggests that transformers, through their strong expressiveness, may exhibit a higher level of algorithms compared to classical human-designed algorithms. Additional analyses and discussions on the transformers in algorithm learning comparing human-designed algorithms can be found in Appendix B.3.
6 Conclusion and Future Works
@@
In this paper, we introduce a novel AlgoFormer, an algorithm learner designed from the looped transformer, distinguished by its algorithmic structures. Comprising three sub-transformers, each playing a distinct role in algorithm learning, the AlgoFormer demonstrates higher expressiveness while maintaining a lower parameter size compared to standard transformers. Moreover, AlgoFormer can represent algorithms more efficiently, in contrast to the vanilla looped transformer. Theoretical analysis establishes that the AlgoFormer can tackle challenging in-context learning tasks, mirroring human-designed algorithms. Our experiments further validate our claim, showing that the proposed transformer outperforms both the standard transformer and the vanilla looped transformer in specific algorithm learning tasks.
There are some potential topics for future research. For example, AlgoFormer exhibits a strikingly similar structure to diffusion models. This similarity arises from the fact that both the AlgoFormer (algorithms) and diffusion models draw inspiration from dynamical systems. Consequently, an interesting question arises: can the empirical training techniques and theoretical insights from the well-explored diffusion model literature be translated to the AlgoFormer for understanding the transformer mechanism?
Appendix A Technical Proof
@@
Notation. We use boldface capital and lowercase letters to denote matrices and vectors respectively. Non-bold letters represent the elements of matrices or vectors, or scalers. For example, $A_{i,j}$ denotes the $(i,j)$-th element of the matrix $\bm{A}$. We use $\|\cdot\|_{2}$ to denote the 2-norm (or the maximal singular value) of a matrix.
@@
According to the construction outlined in Lemma A.2, there exists a one-layer, two-head transformer $\text{TF}_{\text{loop}}$, independent of the input data samples (tokens), that can implement gradient descent for finding the optimal weight $\bm{A}^{*}$ in the context of multivariate regression. The optimization aims to minimize the following empirical loss:
@@
The following lemma highlights the intrinsic “copying” capability of transformers, a pivotal feature for autoregressive models, especially in the context of time series analysis.
@@
According to Lemma 5 in (Li et al., 2023), a seven-layer, two-head pre-transformer $\text{TF}_{\text{pre}}$ is introduced for preprocessing the input CoT sequence. This pre-transformer performs filtering, transforming the input sequence into the structured form given by Equation (7).
@@
Specifically, it identifies the positional index $\ell-1$ of the last token $\hat{\bm{s}}^{\ell-1}$, retains only $\bm{s}_{i}^{\ell-1}$ and $\bm{s}_{i}^{\ell}$ for $1\leq i\leq N$, and filters out all other irrelevant tokens. In this context, the representation function $\Phi^{*}(\cdot)$ corresponds to $L$-layer leaky ReLU MLPs. Notably, the transformation $\bm{s}{i}^{\ell}=\sigma\left(\bm{W}^{\ell}\bm{s}{i}^{\ell-1}\right)$ is expressed, where $\bm{W}^{\ell}$ denotes the weight matrix at the $\ell$-th layer, and $\sigma(\cdot)$ represents the leaky ReLU activation function. Given the reversibility and piecewise linearity of the leaky ReLU activation, we can assume, without loss of generality, that $\bm{s}_{i}^{\ell}=\bm{W}^{\ell}\bm{s}_{i}^{\ell-1}$ in Equation (7). Consequently, the problem is reduced to a multi-variate regression, and a one-layer two-head transformer $\text{TF}_{\text{loop}}$ is demonstrated to effectively implement gradient descent for determining the weight matrix $\bm{W}^{\ell}$, as shown in Lemma A.2. Subsequently, the post-transformer $\text{TF}_{\text{post}}$ produces the desired result $\sigma\left(\bm{W}^{\ell}\hat{\bm{s}}{i}^{\ell-1}\right)$.
@@
In this section, we first show that the one-layer two-head transformer can implement a single step of Newton’s method in Equation (4), with the special form of input token vectors. Then, we introduce the pre-transformer, designed to convert general input tokens into the prescribed format conducive to the transformer’s operation. Finally, the post-transformer facilitates the extraction of the desired results through additional computations, given that the output from the looped-transformer corresponds to an intermediate product.
@@
A transformer with one layer and two heads is capable of implementing one step of Newton’s method in the linear regression problem in Equation (3).
@@
In the next layer, analogous to the construction in Equation (8), we define the following transformations:
@@
The one-layer two-head decoder-based transformer can implement one step of gradient descent in linear regression problems (5).
@@
[**$\left[\begin{array}[]{cccccc}*&*&\cdots&\sum_{j=1}^{i-2}\bm{w}^{i-1\top}_{k}\bm{x}_{j}\bm{x}_{j}&*&\sum_{j=1}^{i-1}\bm{w}^{i\top}_{k}\bm{x}_{j}\bm{x}_{j}\end{array}\right].$
@@
[**$\left[\begin{array}[]{cccccc}*&*&\cdots&-\sum_{j=1}^{i-2}y_{j}\bm{x}_{j}&*&-\sum_{j=1}^{i-1}y_{j}\bm{x}_{j}\end{array}\right]$
@@
Appendix B Experiment
This section provides a comprehensive overview of our experiments, including detailed explanations of our experimental methodology, data/task generation processes, and the selection of hyperparameters. Subsequently, we present extended experiments aimed at exploring the AlgoFormer’s performance under various hyperparameter configurations.
B.1 Experimental Settings and Hyperparameters
Our experiments uniformly utilize 40 in-context samples. The token vectors are represented in $d=20$ dimensional feature vectors, accompanied by $D=256$ dimensional positional embeddings. We choose the Adam optimizer, setting the learning rate $\eta=1e-4$, and employ default parameters for running 500K iterations to achieve convergence. The standard transformer is designed to have $L=12$ layers while pre-, looped and post-transformers are all in one-layer. The default setting for the AlgoFormer, as well as the vanilla looped transformer, involves setting $(T,\Delta T)=(20,15)$.
Sparse linear regression. For the noisy linear regression, as defined in Equations (3) and (5), we start the process by sampling a weight matrix $\bm{w}\in\mathbb{R}^{20}$. Input-label pairs $\{(\bm{x}_{i},y_{i})\}_{i=1}^{40}$ are generated, where $\bm{x}_{i}\sim\mathcal{N}\left(\bm{0},\bm{I}_{20}\right)$, $\epsilon_{i}\sim\mathcal{N}\left(0,1\right)$ and $y_{i}=\bm{w}^{\top}\bm{x}_{i}+\epsilon_{i}$. In the case of sparse linear regression, 85% of the elements in $\bm{w}$ are masked.
Regression with representation. In this task, we instantiate a $3$-layer leaky ReLU MLPs, denoted as $\Phi^{*}(\cdot)$, which remains fixed across all tasks. The data generation process involves sampling a weight matrix $\bm{A}\in\mathbb{R}^{1\times 20}$. Subsequently, input-label pairs $\{(\bm{x}_{i},y_{i})\}_{i=1}^{40}$ are generated, where $\bm{x}_{i}\sim\mathcal{N}\left(\bm{0},\bm{I}_{20}\right)$, $\epsilon_{i}\sim\mathcal{N}\left(0,\sigma^{2}\right)$ and $y_{i}=\bm{A}\Phi^{*}(\bm{x}_{i})+\epsilon_{i}$. In Figure 3a, we specifically set $\sigma=0$. Additionally, we explore the impact of different noise levels by considering $\sigma=0.1$ and $\sigma=1$.
AR(q) with representation. For this task, we set $q=3$ and employ a $3$-layer leaky ReLU MLP denoted as $\Phi^{*}(\cdot)$, consistent across all instances. The representation function accepts a $60$-dimensional vector as input and produces $20$-dimensional feature vectors. The time series sequence $\{\bm{x}_{t}\}_{t=1}^{N}$ is generated by initially sampling $\bm{A}\sim\mathcal{N}\left(\bm{0},\bm{I}_{20\times 20}\right)$. Then the sequence is auto-regressively determined, with $\bm{x}_{t+1}=\bm{A}\Phi^{*}\left([\bm{x}_{t+1-q},\cdots,\bm{x}_{t}]\right)+\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim\mathcal{N}\left(\bm{0},\bm{I}_{20}\right)$.
CoT with MLPs. In this example, we generate a 6-layer leaky ReLU MLP to serve as a CoT sequence generator, determining the length of CoT steps for each sample to be six. The CoT sequence, denoted as $\{\bm{x},\bm{s}^{1},\cdots,\bm{s}^{L}\}$ is generated by first sampling $\bm{x}\sim\mathcal{N}\left(\bm{0},\bm{I}_{20}\right)$, where $\bm{s}^{\ell}\in\mathbb{R}^{20}$ represents the intermediate state output from the $\ell$-th layer of the MLP.
Linear regression with gradient descent and Newton’s method. We compare the AlgoFormer with the gradient descent and Newton’s method in solving classical linear regression problems. The experiments are conducted with varying noise levels, specifically $\sigma\in\{0.0,0.1,1.0\}$. The hyperparameters for AlgoFormer are kept at their default settings, as mentioned earlier. For the gradient descent, we perform a grid search for the step size, exploring values from $\{1e0,5e-1,1e-1,5e-2,1e-2,1e-3\}$. The initialization scaler $\alpha$ is selected from $\{\frac{1}{4}\alpha_{\max},\frac{1}{2}\alpha_{\max},\frac{3}{4}\alpha_{\max}\}$, where $\alpha_{\max}:=\frac{2}{\left\|\bm{S}\bm{S}^{\top}\right\|_{2}}$ in Equation (4). The reported results represent the best outcomes obtained through hyperparameter grid search.
B.2 Impact of Hyperparameters
(a) $\Delta T=5$
(b) $\Delta T=10$
(c) $\Delta T=15$
Figure 4: The validation error of trained models, evaluated on regression with representation task, with varying hyperparameters $T$ and $\Delta T$. The AlgoFormers are trained for $T$ loops, defined in Equation (6), and the evaluation focuses on square loss at longer iterations, where the number of loop iterations far exceeds $T$.
(a) Varying numbers of layers
(b) Varying numbers of heads
Figure 5: The validation error of trained models, evaluated on regression with representation task, with varying numbers of layers (denoted as $L$) and heads (denoted as $h$). In the context of AlgoFormer, the number of layers $L$ corresponds to the layers in the pre-, looped, and post-transformers, all of which are $L$-layer transformers. The AlgoFormers are trained with $(T,\Delta T)=(20,15)$, defined in Equation (6).
(a) Linear regression, $\sigma=0$
(b) Linear regression, $\sigma=0.1$
(c) Linear regression, $\sigma=1.0$
Figure 6: The validation error of trained AlgoFormer models and the linear regression models optimized by gradient descent and Newton’s method (with different iteration steps). The AlgoFormers are trained with $(T,\Delta T)=(20,15)$ and $(T,\Delta T)=(10,10)$, defined in Equation (6). Here, Newton and GD denote the Newton’s method in Equation (4) and classical gradient descent, respectively.
Loop iterations. We conduct comprehensive experiments on the AlgoFormer with varying loop numbers, on solving the regression with representation task. The results highlight the crucial role of both $T$ and $\Delta T$ in the performance of the AlgoFormer. It is observed that a larger $\Delta T$ contributes to the stable inference of transformers. Comparing Figure 4a with Figures 4b and 4c, it is evident that a larger $\Delta T$ enhances stable long-term inference. The number of loop iterations $T$ determines the model capacity in expressing algorithms. However, it is important to note that there exists a trade-off between the iteration numbers $(T,\Delta T)$ and computational costs. Larger $(T,\Delta T)$ certainly increases model capacity but also leads to higher computational costs and challenges in model training, as reflected in Figure 4.
Number of heads and layers. In our experiments on the AlgoFormer, we vary the numbers of heads and layers in $\text{TF}_{\text{loop}}$ while addressing the regression with representation task. The results reveal a consistent trend that an increase in both the number of heads and layers leads to lower errors. This aligns with our intuitive understanding, as transformers with greater numbers of heads and layers exhibit enhanced expressiveness. However, a noteworthy observation is that 4-layer and 16-head transformers may exhibit suboptimal performance, possibly due to increased optimization challenges during model training. This finding underscores the importance of carefully selecting the model size, as a larger model, while offering higher expressiveness, may present additional training difficulties. The visualized results are shown in Figure 5. Moreover, compared with the standard transformer (GPT-2), even with the same number of layers, the AlgoFormer exhibits better performance, mainly due to the introduced algorithmic structure. This finding highlights the role of the regularization of model structure. Therefore, we have reasons to believe that the good performance of the AlgoFormer not only comes from the higher expressiveness with deeper layers but also from the regularization of model architecture, which facilitates easier training and good generalization.
B.3 Comparison with Newton’s Method and Gradient Descent
In this subsection, we compare the AlgoFormer with Newton’s method and gradient descent in solving linear regression problems. We adopt the same default hyperparameters as specified in Appendix B.1.
Several aspects contribute to this phenomenon. Firstly, in the noiseless case, Newton’s method can precisely recover the weights through the linear regression objective in Equation (5), capitalizing on its superlinear convergence. On the other hand, the AlgoFormer operates as a black-box, trained from finite data. While we demonstrate good model expressiveness, the final generalization error of the trained transformer results from the model’s expressiveness, the finite number of training samples, and the optimization error. Despite exhibiting high expressiveness, the trained AlgoFormer cannot eliminate the last two errors entirely. This observation resonates with similar findings in solving partial differential equations (Raissi et al., 2019) and large linear systems (Gu & Ng, 2023) using deep learning models.
Secondly, with larger noise levels, Newton’s method shows suboptimal results. This is partly due to the inclusion of noise, which slows down the convergence rate, and Newton’s method experiences convergence challenges when moving away from the local solution. In terms of global convergence, the transformer demonstrates superior performance compared to Newton’s method.
In conclusion, although the transformer may be smarter than human-designed algorithms (Newton’s method and gradient descent), it currently cannot replace these classical algorithms in specific scientific computing tasks. Human-designed algorithms, backed by problem priors and precise computation, achieve irreplaceable performance. It’s important to note that deep learning models, including transformers, are specifically designed for solving black-box tasks where there is limited prior knowledge but sufficient observation samples. We expect that transformers, with their substantial expressiveness, hold the potential to contribute to designing implicit algorithms in solving scientific computing tasks.

### v2

@@ 5.3 Impact of Hyperparameters
In this subsection, we conduct the empirical analysis of the impact of the hyperparameters on the AlgoFormer.
(a) $\Delta T=5$
(b) $\Delta T=10$
(c) $\Delta T=15$
Figure 3: The validation error of trained models, evaluated on regression with representation task, with varying hyperparameters $T$ and $\Delta T$. The AlgoFormers are trained for $T$ loops, defined in Equation 3, and the evaluation focuses on square loss at longer iterations, where the number of loop iterations far exceeds $T$.
(a) Varying numbers of layers
(b) Varying numbers of heads
Figure 4: The validation error of trained models, evaluated on regression with representation task, with varying numbers of layers (denoted as $L$) and heads (denoted as $h$). In the context of AlgoFormer, the number of layers $L$ corresponds to the layers in the pre-, looped, and post-transformers, all of which are $L$-layer transformers. The AlgoFormers are trained with $(T,\Delta T)=(20,15)$, defined in Equation 3.
(a) Linear regression, $\sigma=0$
(b) Linear regression, $\sigma=0.1$
(c) Linear regression, $\sigma=1.0$
Figure 5: The validation error of trained AlgoFormer models and the linear regression models optimized by gradient descent and Newton’s method. The AlgoFormers are trained with $(T,\Delta T)=(20,15)$ and $(T,\Delta T)=(10,10)$, defined in Equation 3.
Loop iterations. We conduct comprehensive experiments on the AlgoFormer with varying loop numbers, on solving the regression with representation task. The results highlight the crucial role of both $T$ and $\Delta T$ in the performance of the AlgoFormer. It is observed that a larger $\Delta T$ contributes to the stable inference of transformers. Comparing Figure 3a with Figures 3b and 3c, it is evident that a larger $\Delta T$ enhances stable long-term inference. The number of loop iterations $T$ determines the model capacity in expressing algorithms. However, it is important to note that there exists a trade-off between the iteration numbers $(T,\Delta T)$ and computational costs. Larger $(T,\Delta T)$ certainly increases model capacity but also leads to higher computational costs and challenges in model training, as reflected in Figure 3.
Number of heads and layers. In our experiments on the AlgoFormer, we vary the numbers of heads and layers in $\text{TF}_{\text{loop}}$ while addressing the regression with representation task. The results reveal a consistent trend that an increase in both the number of heads and layers leads to lower errors. This aligns with our intuitive understanding, as transformers with greater numbers of heads and layers exhibit enhanced expressiveness. However, a noteworthy observation is that 4-layer and 16-head transformers may exhibit suboptimal performance, possibly due to increased optimization challenges during model training. This finding underscores the importance of carefully selecting the model size, as a larger model, while offering higher expressiveness, may present additional training difficulties. The visualized results are shown in Figure 4. Moreover, compared with the standard transformer, even with the same number of layers, the AlgoFormer exhibits better performance in those tasks, mainly due to the introduced algorithmic structure. This finding highlights the role of the structure regularization of the model. Therefore, we have reasons to believe that the good performance of the AlgoFormer not only comes from the higher expressiveness with deeper layers but also from the structure regularization of model architecture, which facilitates easier training and good generalization.
In Figure 3, we also provide insights into the behavior of AlgoFormer during inference. Although the AlgoFormer is trained with T=20 loop iterations, the inference is conducted with significantly longer loop iterations (e.g., T=200), As T increases beyond the training length, AlgoFormer demonstrates improved performance and eventually stabilizes (i.e., converges). This behavior suggests that the inner looped transformers are effectively performing iterative computations that converge to a stable solution.
5.4 AlgoFormer and Human-Designed Algorithms
In this subsection, we compare the AlgoFormer with Newton’s method and gradient descent in solving linear regression problems. We adopt the same default hyperparameters, with their selection grounded in a comprehensive grid search.
Several aspects contribute to this phenomenon. Firstly, in the noiseless case, Newton’s method can precisely recover the weights through the linear regression objective in Equation 6, capitalizing on its superlinear convergence. On the other hand, the AlgoFormer operates as a black-box, trained from finite data. While we demonstrate good model expressiveness, the final generalization error of the trained transformer results from the model’s expressiveness, the finite number of training samples, and the optimization error. Despite exhibiting high expressiveness, the trained AlgoFormer cannot eliminate the last two errors entirely. This observation resonates with similar findings in solving partial differential equations (Raissi et al., 2019) and large linear systems (Gu & Ng, 2023) using deep learning models.
Secondly, with larger noise levels, Newton’s method shows suboptimal results. This is partly due to the inclusion of noise, which slows down the convergence rate, and Newton’s method experiences convergence challenges when moving away from the local solution. In terms of global convergence, the AlgoFormer demonstrates superior performance compared to Newton’s method.
Human-designed algorithms, backed by problem priors and precise computation, achieve irreplaceable performance. It’s important to note that deep learning models, including transformers, are specifically designed for solving black-box tasks where there is limited prior knowledge but sufficient observation samples. We expect that transformers, with their substantial expressiveness, hold the potential to contribute to designing implicit algorithms in solving scientific computing tasks.
5.5 Applications to Language Tasks
In this subsection, we extend the evaluation of the proposed AlgoFormer to real-world language tasks, complementing its performance in real applications. Specifically, we focus on Neural Machine Translation using the IWSLT 2015 German-English dataset. The experimental setup includes a standard Transformer with 12 layers, 8 attention heads, a feature dimension of 256, and a learning rate of 5e-5. The pre-, looped, and post-transformers are all implemented as single layers, with $T$ set to 10 and $\Delta T$ to 10 (see the hyperparameters in Equation 3 for training). The input German text is treated as a prefix, and the output English text is generated autoregressively using decoder-based transformers for all three models. The translation performance is evaluated using cross-entropy loss, where lower values indicate better results. Additionally, BLEU (Bilingual Evaluation Understudy) is a widely used metric for measuring the quality of translation tasks, with higher scores indicating better performance. The results are presented in the table below:
Model Standard Looped AlgoFormer
Cross entropy 4.99 4.73 4.61
BLUE 9.30 10.56 14.72
Table 1: Quality of Transformer models on machine translation.
We also implement the proposed AlgoFormer on the text classification task using various datasets (AG News, IMDB, DBPedia, Yelp Review, and Yahoo News) to evaluate its performance on a real-world language application. The experimental setup includes a standard Transformer with 12 layers, 1 attention head, a feature dimension of 32, and a learning rate of 1e-3. The pre-, looped, and post-transformers are all implemented as single layers with one attention head, with $(T,\Delta T)=(10,10)$. The input news text data is processed using encoder-based transformers, and the output classification is generated. Classification accuracy, where higher accuracy indicates better performance, is used as the evaluation metric. The results are summarized in the table below:
Model Standard Looped AlgoFormer
AG News 92.07 91.04 97.92
IMDB 75.00 77.50 77.50
DBPedia 99.21 99.11 98.21
Yelp Review 66.25 53.57 57.50
Yahoo News 73.96 69.79 81.25
Table 2: Accuracy (%) of Transformer models on text classification across different datasets.
As shown in the results, AlgoFormer outperforms and performs comparably with the standard Transformer and the vanilla looped Transformer in the Neural Machine Translation and Text Classification tasks, suggesting that conventional models may have inherent redundancies. This aligns with recent findings on the redundancy in large language models (Chen et al., 2024; Frantar & Alistarh, 2023; Xia et al., 2024). These preliminary results indicate the potential of AlgoFormer, which is designed as an algorithmic conductor, in real-world language tasks.
5.6 Computation Comparison
The additional computational costs and complexity introduced by AlgoFormer over the standard transformer and the vanilla looped transformer depend on the specific structure of the AlgoFormer framework. In our experimental settings, these costs are negligible. Compared to the standard transformer, the total computation overhead is primarily determined by the loop iterations $T$ and the number of loops $\Delta T$ included in the training loss. Since $T$ in our experiments is comparable to the number of layers in the standard transformer, the dominant additional computation comes from $\Delta T$. This overhead can be mitigated by using smaller batch sizes, ensuring that the proposed method does not introduce significant computational overhead compared to the standard transformer. The main complexity lies in the training strategy outlined in Equation 3, which ensures stable training.
When compared to the vanilla looped transformer, the computational overhead comes from the pre- and post-transformer modules. However, if these modules are of similar size to the looped transformer, the additional computational cost is negligible, especially when the loop iteration $T$ is much greater than the number of layers in these modules. For instance, in our experiments, we set the pre-, looped-, and post-transformer layers to 1, with $T=20$ during training. In this scenario, the additional computation introduced by the pre- and post-transformer modules constitutes approximately 1/10 of the total computation. During inference, where $T$ is typically much larger than during training (e.g., T=200 loop iterations in Figure 3), the additional computational overhead is further reduced to about 1/100.
6 Conclusion and Discussion
@@
In this paper, we introduce a novel AlgoFormer, an algorithm learner designed from the looped transformer, distinguished by its algorithmic structures. We provide an insight that the efficient transformer framework can be designed by considering the prior knowledge of the task and the structure of potential algorithms. Comprising three sub-transformers, each playing a distinct role in algorithm learning, the AlgoFormer demonstrates expressiveness and efficiency while maintaining a low parameter size. Theoretical analysis establishes that the AlgoFormer can tackle challenging in-context learning tasks, mirroring human-designed algorithms. Our experiments further validate our claim, showing that the proposed transformer outperforms both the standard transformer and the vanilla looped transformer in specific algorithm learning and real-world language tasks.
While the proposed AlgoFormer framework offers flexibility and efficiency, several potential limitations and challenges warrant discussion.
First, the dependence on prior knowledge for designing concrete architecture. The design of the AlgoFormer framework relies heavily on prior knowledge of tasks and the structure of potential algorithms. In scientific computing tasks, where problems are often well-formulated and informed by domain knowledge, this reliance can be an advantage, enabling the design of task-specific, efficient architectures. However, for real-world language tasks, the abstract and unstructured nature of the problems often limits the availability of such prior knowledge. This can make it challenging to design an optimal AlgoFormer architecture for these tasks. While the insights of AlgoFormer can significantly benefit the scientific computing community, extending its design principles to language tasks may require more advanced techniques, such as automated architecture search or meta-learning, to determine suitable architectures without extensive prior knowledge.
Second, the training complexity of AlgoFormer with more complicated algorithmic structures. Although incorporating more complex algorithmic structures into AlgoFormer, such as nested looped transformers, enhances its expressiveness, training AlgoFormer can become more challenging. It inevitably introduces additional computational costs and requires adjustments to the training strategy. These challenges could limit the usability of AlgoFormer in scenarios where algorithmic structures are complicated to model.
Third, the scalability of AlgoFormer to large-scale applications. The scaling behavior of AlgoFormer has not been thoroughly investigated in this paper. The tasks studied are primarily hand-generated scientific problems or toy cases, and the language task considered is of medium scale. Similarly, the model sizes used in this study are relatively small compared to state-of-the-art large language models (LLMs). To fully assess the potential of AlgoFormer, future research should explore its scalability to larger, more challenging tasks and datasets, as well as the associated computational and memory requirements when deploying AlgoFormer at scale.
Acknowledgements
M. Ng’s research is partly supported by the National Key Research and Development Program of China under Grant 2024YFE0202900; HKRGC GRF 17201020 and 17300021, HKRGC CRF C7004-21GF, and Joint NSFC and RGC N-HKU769/21.
Appendix A Technical Proofs
@@
Notation. We use boldface capital and lowercase letters to denote matrices and vectors respectively. Non-bold letters represent the elements of matrices or vectors, or scalars. For example, $A_{i,j}$ denotes the $(i,j)$-th element of the matrix $\bm{A}$. We use $\|\cdot\|_{2}$ to denote the 2-norm (or the maximal singular value) of a matrix.
@@
According to the construction outlined in Lemma A.2, there exists a one-layer, two-head transformer $\text{TF}_{\text{loop}}$, independent of the exact values of input data samples (tokens), that can implement gradient descent for finding the optimal weight $\bm{A}^{*}$ in the context of multivariate regression. The optimization aims to minimize the following empirical loss:
@@
The following lemma highlights the intrinsic “copying" capability of transformers, a pivotal feature for autoregressive models, especially in the context of time series analysis.
@@
According to Lemma 5 in (Li et al., 2023), a seven-layer, two-head pre-transformer $\text{TF}_{\text{pre}}$ is introduced for preprocessing the input CoT sequence. This pre-transformer performs filtering, transforming the input sequence into the structured form given by Equation 7.
@@
Specifically, it identifies the positional index $\ell-1$ of the last token $\hat{\bm{s}}^{\ell-1}$, retains only $\bm{s}_{i}^{\ell-1}$ and $\bm{s}_{i}^{\ell}$ for $1\leq i\leq N$, and filters out all other irrelevant tokens. In this context, the representation function $\Phi^{*}(\cdot)$ corresponds to $L$-layer leaky ReLU MLPs. Notably, the transformation $\bm{s}{i}^{\ell}=\sigma\left(\bm{W}^{\ell}\bm{s}{i}^{\ell-1}\right)$ is expressed, where $\bm{W}^{\ell}$ denotes the weight matrix at the $\ell$-th layer, and $\sigma(\cdot)$ represents the leaky ReLU activation function. Given the reversibility and piecewise linearity of the leaky ReLU activation, we can assume, without loss of generality, that $\bm{s}_{i}^{\ell}=\bm{W}^{\ell}\bm{s}_{i}^{\ell-1}$ in Equation 7. Consequently, the problem is reduced to a multi-variate regression, and a one-layer two-head transformer $\text{TF}_{\text{loop}}$ is demonstrated to effectively implement gradient descent for determining the weight matrix $\bm{W}^{\ell}$, as shown in Lemma A.2. Subsequently, the post-transformer $\text{TF}_{\text{post}}$ produces the desired result $\sigma\left(\bm{W}^{\ell}\hat{\bm{s}}{i}^{\ell-1}\right)$.
@@
In this section, we first show that the one-layer two-head transformer can implement a single step of Newton’s method in Equation 5, with the special form of input token vectors. Then, we introduce the pre-transformer, designed to convert general input tokens into the prescribed format conducive to the transformer’s operation. Finally, the post-transformer facilitates the extraction of the desired results through additional computations, given that the output from the looped-transformer corresponds to an intermediate product.
@@
A transformer with one layer and two heads is capable of implementing one step of Newton’s method in the linear regression problem in Equation 4.
@@
In the next layer, analogous to the construction in Equation 8, we define the following transformations:
@@
The one-layer two-head decoder-based transformer can implement one step of gradient descent in linear regression problems in Equation 6.
@@
$\left[\begin{array}[]{cccccc}*&*&\cdots&\sum_{j=1}^{i-2}\bm{w}^{i-1\top}_{k}\bm{x}_{j}\bm{x}_{j}&*&\sum_{j=1}^{i-1}\bm{w}^{i\top}_{k}\bm{x}_{j}\bm{x}_{j}\end{array}\right].$
@@
$\left[\begin{array}[]{cccccc}*&*&\cdots&-\sum_{j=1}^{i-2}y_{j}\bm{x}_{j}&*&-\sum_{j=1}^{i-1}y_{j}\bm{x}_{j}\end{array}\right]$